OpenStudy (anonymous):

solve for x and show steps: 4^(3x)=1/√2

OpenStudy (anonymous):

(2^2)^(3x) = 2^(-1/2)

OpenStudy (anonymous):

6x=-1/2

OpenStudy (anonymous):

x=-1/12

OpenStudy (anonymous):

@raheen where did the "-" came from >??????

OpenStudy (anonymous):

1/sqrt(2) = 1/2^(1/2) = 2^(-1/2)

OpenStudy (anonymous):

1/a^m = a^(-m)

OpenStudy (anonymous):

okay i got them to the same base where the equation is \[(2^2)^3x=2^-1/2\] but you cant multiply the 2 and 3x because the 3x is a power of the 2

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

its equal to -1/2 sorry